# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v1

@@ 4. Proofs of the main results
In this section we first derive Theorems A and B from Theorem E in Section 4.1. Then, in Section 4.2, we establish the latter as a consequence of Theorem C. In Section 4.3, we turn to the proofs of Theorem C, which is the central part of the argument, and Corollary D. Finally, Corollary F is proved in Section 4.4
@@ 4.1.1. Proof of Part (M1) of Theorem A
When $M$ is convex, $x$ is in fact a least period closed orbit on $M$, among all (not necessarily ${\mathbb{Q}}$-visible) closed orbits. In other words, at least one of the least period closed orbits is necessarily ${\mathbb{Q}}$-visible. This is a consequence of the fact that for convex hypersurfaces the (non-equivariant) homological capacity is equal to the least period, [AbK, Ir].
@@
By (O2), the orbits $x_{j}$, $j\geq 0$, form one cluster, i.e., have the same ratio ${\mathcal{A}}(x_{j})/\operatorname{\hat{\mu}}(x_{j})$. Without loss of generality, we may assume that this ratio is 1, i.e.,
@@
We already know that both sides in (D1) and (D2) are supported in one degree and one-dimensional. Hence we only need to determine the degree shift $L$. Let $\varphi$ be the Poincaré return map of $x$. Then the return map of $x^{l}$ is $\varphi^{l}$. By (2.10) and (2.16),
@@ 4.2. Proof of Theorem E
When the flow is non-degenerate, Part (O1) has an analogue for any field ${\mathbb{F}}$ with $\deg(x)$ replaced my $\mu(x)$, although the statement is then less precise. Namely, the sequences are $\mu(x_{i})$ and $\deg(I_{1})$ are still increasing although now not necessarily strictly. However, for every $m\in n-1+2{\mathbb{N}}$ we still have exactly one ${\mathbb{F}}$-visible closed orbit $x$ with $\mu(x)=m$ and all $\deg(I_{i})$ have parity $n$. The key difference is that in this case we cannot rule out the existence of closed ${\mathbb{F}}$-visible orbits $x$ with $\mu(x)$ of parity $n$ and intervals in the sequences $\deg(x_{i})$ and $\deg(I_{i})$ where these sequences are constant. The proof is a subset of the proof of Part (O1) combined with the fact that $\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})=\{\mu(x),\mu(x)+1\}$ for any field ${\mathbb{F}}$ when $x$ is non-degenerate, i.e., the second part of Corollary D holds automatically. We do not know to what degree Part (O2) extends to this case.
@@ 4.3.1. Generalities
As above, by the universal coefficient theorem and, in particular (2.4), it is enough to prove (1.1) for all postive characteristic fields ${\mathbb{F}}_{p}$. Setting
@@
$\mu_{+}\big{(}x_{ij}^{k}\big{)}\leq d_{il}-2\textrm{ when }{\mathcal{A}}\big{(}x_{ij}^{k}\big{)}<C_{l}$ (4.13)
@@
Here both action inequalities are strict, since the sequence $C_{l}$ can be chosen outside the action spectrum ${\mathcal{S}}(\alpha)$.
@@
(Here we are using the fact that $i=i^{\prime}$, i.e., the orbits are in the same cluster.) Due to (B2), we have equality in both of these inequalities and the index difference is exactly $2$. Therefore, by (B3), $\deg(I)=\mu_{-}\big{(}x_{ij}^{k}\big{)}+1=d_{il}+n$. ∎
@@
At the same time, the left-hand side is equal $(-1)^{n}D$ by Lemma 4.9. Indeed, each term on the left is either 00 when $m\neq d_{il}+n$ or $(-1)^{n}\dim\operatorname{SH}^{C_{l}}_{m}(W)$ when $m=d_{il}+m$ since $d_{il}$ is even. Thus

### v2

@@ 4. Proofs of the main results
Here we first derive Theorems A and B from Theorem E in Section 4.1. Then, in Section 4.2, we establish the latter as a consequence of Theorem C. In Section 4.3, we turn to the proofs of Theorem C, which is the central part of the argument, and Corollary D. Finally, Corollary F is proved in Section 4.4
@@ 4.1.1. Proof of Part (M1) of Theorem A
When $M$ is convex, $x$ is, in fact, a least period closed orbit on $M$, among all (not necessarily ${\mathbb{Q}}$-visible) closed orbits. In other words, at least one of the least period closed orbits is necessarily ${\mathbb{Q}}$-visible. This is a consequence of the fact that for convex hypersurfaces the (non-equivariant) homological capacity is equal to the least period, [AbK, Ir].
@@
By (O2) and homogeneity of ${\mathcal{A}}$ and $\operatorname{\hat{\mu}}$, the orbits $x_{j}$, $j\geq 0$, form one cluster, i.e., have the same ratio ${\mathcal{A}}(x_{j})/\operatorname{\hat{\mu}}(x_{j})$. Without loss of generality, we may assume that this ratio is 1, i.e.,
@@
We already know that $\operatorname{CH}(x;{\mathbb{Q}})$ and $\operatorname{CH}\big{(}x^{k};{\mathbb{Q}}\big{)}$ are supported in one degree and one-dimensional. Hence we only need to determine the degree shift $L$. Let $\varphi$ be the Poincaré return map of $x$. Then the return map of $x^{l}$ is $\varphi^{l}$. By (2.10) and (2.16),
@@ 4.2. Proof of Theorem E
When the flow is non-degenerate, Part (O1) has an analogue for any field ${\mathbb{F}}$ with $\deg(x)$ replaced my $\mu(x)$, although the statement is then less precise. Namely, the sequences are $\mu(x_{i})$ and $\deg(I_{i})$ are still increasing although now not necessarily strictly. However, for every $m\in n-1+2{\mathbb{N}}$ we still have exactly one ${\mathbb{F}}$-visible closed orbit $x$ with $\mu(x)=m$ and all $\deg(I_{i})$ have parity $n$. The key difference is that in this case we cannot rule out the existence of closed ${\mathbb{F}}$-visible orbits $x$ with $\mu(x)$ of parity $n$ and intervals in the sequences $\deg(x_{i})$ and $\deg(I_{i})$ where these sequences are constant. The proof is a subset of the proof of Part (O1) combined with the fact that $\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})=\{\mu(x),\mu(x)+1\}$ for any field ${\mathbb{F}}$ when $x$ is non-degenerate, i.e., the second part of Corollary D holds automatically. We do not know to what degree Part (O2) extends to this case.
@@ 4.3.1. Generalities
As above, by the universal coefficient theorem and, in particular (2.4), it is enough to prove (1.1) for all positive characteristic fields ${\mathbb{F}}_{p}$. Setting
@@
$\mu_{+}\big{(}x_{ij}^{k}\big{)}\leq d_{il}+n-1\textrm{ when }{\mathcal{A}}\big{(}x_{ij}^{k}\big{)}<C_{l}$ (4.13)
@@
Indeed, in the setting of Theorem 3.3 and Corollary 3.5 we have $k\leq k_{ijl}$ in (4.13) and $k>k_{ijl}$ in (4.14). Here both action inequalities are strict, since the sequence $C_{l}$ can be chosen outside the action spectrum ${\mathcal{S}}(\alpha)$.
@@
(Here we are using the fact that $i=i^{\prime}$, i.e., the orbits are in the same cluster.) Due to (B2), we have equality in both of these inequalities and the index difference is exactly $2$. Therefore, by (B3), $\deg(I)=\mu_{+}\big{(}x_{ij}^{k}\big{)}+1=d_{il}+n$. ∎
@@
At the same time, the left-hand side is equal $(-1)^{n}D$ by Lemma 4.9. Indeed, each term on the left is either 00 when $m\neq d_{il}+n$ or $(-1)^{n}\dim\operatorname{SH}^{C_{l}}_{m}(W)$ when $m=d_{il}+n$ since $d_{il}$ is even. Thus